OpenStudy (anonymous):

evaluate f"(c) at the point. f(x) = 3-x/2x+3 C=-2

OpenStudy (anonymous):

f(x)=(3-x)/(2x+3), u = 3-x , v= 2x+3 , u'=-1 , v'=2 f'(x) = u'v+uv' -2x-3+ 6-2x ------- = ------------ , u = -2x-3+ 6-2x , v = (2x+3)^2 v^2 (2x+3)^2 f''(x) = u'v+uv' , use the same way ------- v^2

OpenStudy (anonymous):

ok i was doing it the right way yah:) lol

OpenStudy (anonymous):

ok g'luck

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

wait, a sec, this looks like a quotient rule problem, yet u state the product rule?

OpenStudy (anonymous):

yeah i think i'm lost at the f''